How would you solve the equation log2(x)^2=(log2(x))^2 I am stuck on how to square a log
Answer by stanbon(75887)

How would you solve the equation
log2(x)^2 = (log2(x))^2
--------------
2log2(x) = [log2(x)]^2
----
Rearrange:
[log2(x)]^2 - 2*log2(x) = 0
----
Factor:
log2(x)[log2(x)-2] = 0
----
Solve for "x":
log2(x) = 0 or log2(x)-2 = 0
---
x = 2^0 or log2(x) = 2
---
x = 1 or x = 2^2 = 4
---
x = 1 or x = 4
=====================
Cheers,
Stan H.
=====================